OpenStudy (anonymous):

Which of the following shows the equation below correctly solved for m? mn = 3m − 2 A. m=3m-2/n B. m=-2/n-3 C. m=mn+2/3 D. m=n+2/3

OpenStudy (anonymous):

Sorry, I didn't see that other m

OpenStudy (anonymous):

mn = 3m − 2 -3m ----------------- mn - 3m = -2 m(n-3) = -2 m = -2/ n-3

OpenStudy (anonymous):

B is the answer